NicktheBlade16:

What is the simplified form of 6 log3x - 7 log3y + 4 log3z? log base 3 of the quantity negative xz end quantity divided by 4y end log log base 3 of 3xz divided by y end log log base 3 of the quantity x to the sixth z to the fourth end quantity divided by y to the seventh end log log base 3 of the quantity xy all raised to the twelfth end quantity divided by z to the 4th end log

NicktheBlade16:

Alexis1415:

The answer would be, log base 3 of the quantity x to the seventh z to the fourth end quantity divided by y to the sixth end log \[6\log _{3}x-7\log _{3}y+4\log _{3}z\] You have to use the power law, \[alog _{x}b=\log _{x}b ^{a}\] so \[\log _{3}x ^{6}-\log _{3}y ^{7}+\log _{3}z ^{4}\] With all of this it shows it's the third option c:

Alexis1415:

I didn't realize how spaced out I put it whoops

NicktheBlade16:

Its perfectly fine.

Arieonna:

i would say C

Alexis1415:

yeah thats what i said :0

NicktheBlade16:

yeah

Arieonna:

wait u did